OpenStudy (anonymous):

Multiply Matrices Help!!! Find -5A+4B A=[6 1 over -4 -6 over 7 -7] B=[-5 -1 over -3 -8 over 6 8]

OpenStudy (amistre64):

thats addition of matrixes ... just multiply the scalar to all the elements .. then add the reslts

OpenStudy (amistre64):

A=[6 1 over -4 -6 over 7 -7] B=[-5 -1 over -3 -8 over 6 8] -5A=[ 6(-5) 1(-5) over -4(-5) -6(-5) over 7(-5) -7(-5)] +4B=[-5(4) -1(4) over -3(4) -8(4) over 6(4) 8(4) ] ---------------------------------------------------- add em up

OpenStudy (amistre64):

thats a very nice way of presenting the matrixes ... makes lining up the elements simple lol

OpenStudy (anonymous):

I got 1 -4 over -9 -11 over 2 -12 on the first one and -20 -4 over -12 -32 over 24 32. What do you mean by add them?